# A Survey on Neural Question Generation: Methods, Applications, and Prospects

### v1

@@ 4.2 VQG
VQG can be regarded as a dual task of visual question answering. As illustrated in Figure 1, VQG models are typically divided into three main categories, including CNN-based, GNN-based, and Generative Networks-based.
@@
GNN-based. One key challenge of VQG is to focus on answer-related regions during question generation. To solve the issue, researchers propose several approaches to perform explicit region selection. For instance,  Xu et al. Xu et al. (2020) perform explicit object-level cross-modal interaction by identifying a core answer area and constructing an answer-related graph convolutional network (GCN) graph structure.  Xie et al. Xie et al. (2021) leverage a co-attention network and a graph network to identify and relate key objects in an image to a target answer, thereby generating more comprehensive questions. However, previous approaches rely solely on semantic features to identify regions related to the answer, leading to potential biases and overlooking complex relations between objects. Given this,  Chai et al. Chai et al. (2023) utilize contrastive learning to integrate semantic knowledge with regional representations and leverage a relation-level interaction scenario to consider multiple types of relations among regions and answers.
@@ 5 Hybrid Neural Question Generation
Hybrid NQG aims to generate textual questions based on both structured and unstructured data sources, where multimodal is common. Compared with the previous single-modal question generation, hybrid question generation is more prevalent in real-life scenarios, especially in the education field. The primary challenge in hybrid question generation lies in effectively integrating information across diverse data sources or modalities. As pioneers in the field, Wang and Baraniuk Wang and Baraniuk (2023) first investigate multi-modal question generation from images and texts, proposing a novel and effective PLMs-based approach that surpasses the performance of ChatGPT. Despite the promising results achieved by Wang and Baraniuk Wang and Baraniuk (2023), hybrid NQG remains under-explored, with significant potential for further innovation and improvement.
@@ 6 Conclusion and Future Directions
This paper provides a thorough overview of NQG in various modalities. We first introduce popular datasets, classical evaluation metrics, and four prominent applications. We then explore prevalent methods for modeling diverse inputs. Despite the notable achievements of NQG models, several challenges remain, suggesting potential directions for future research.
@@
Multi-modal Question Generation. Current question generation tasks primarily concentrate on single-modal question generation, such as KBQG, TQG, and VQG. Nevertheless, in practical scenarios, the significance of multi-modal question generation is on the rise. This is particularly evident in the educational field, where numerous scientific questions require an understanding of both visual images and textual descriptions. To the best of our knowledge, multi-modal question generation remains in a very early stage, with few studies having been conducted in this field Wang and Baraniuk (2023), as detailed in Section 5. Given the advanced capabilities of vision-language pre-trained models like CLIP, a compelling research direction is to develop effective strategies for leveraging VL-PTMs in multi-modal question generation.

### v2

@@ 4.2 VQG
VQG may be considered a dual task of visual question answering. As illustrated in Figure 1, VQG models are typically divided into three main categories, including CNN-based, GNN-based, and Generative Networks-based.
@@
GNN-based. One key challenge of VQG is to focus on answer-related regions during question generation. To solve the issue, researchers propose several approaches to perform explicit region selection. For instance,  Xu et al. Xu et al. (2020) perform explicit object-level cross-modal interaction by identifying a core answer area and constructing an answer-related graph convolutional network (GCN) graph structure.  Xie et al. Xie et al. (2021) leverage a co-attention network and a graph network to identify and relate key objects in an image to a target answer, thereby generating more comprehensive questions. However, previous approaches rely solely on semantic features to identify regions related to the answer, leading to potential biases and overlooking complex relations between objects. Given this,  Chai et al. Chai et al. (2023) utilize contrastive learning to integrate semantic knowledge with regional representations and leverage a relation-level interaction scenario to consider various types of relationships between regions and answers.
@@ 5 Hybrid Neural Question Generation
Hybrid NQG aims to generate textual questions based on both structured and unstructured data sources Wang and Baraniuk (2023); Mi et al. (2024); Dong et al. (2024), where multimodal is common. Compared with the previous single-modal question generation, hybrid question generation is more prevalent in real-life scenarios, especially in the education field. The primary challenge in hybrid question generation lies in effectively integrating information across diverse data sources or modalities. As pioneers in the field, Wang and Baraniuk Wang and Baraniuk (2023) first investigate multi-modal question generation from images and texts, proposing a novel and effective PLMs-based approach that surpasses the performance of ChatGPT. Dong et al. Dong et al. (2024) propose a unified framework for generating contextual questions (CQG) and factoid questions (FQG), which addresses current methods’ limitations in structural and contextual information. Specifically, they introduce shared task modules for cross-domain learning and task-specific modules that integrate external knowledge for CQG and enhance contextual understanding for FQG, demonstrating advanced performance. Despite these promising results, hybrid NQG remains under-explored, with significant potential for further innovation and improvement.
@@ 6 Conclusion and Future Directions
This paper provides a thorough overview of Neural Question Generation (NQG) in various modalities. We first introduce popular datasets, classical evaluation metrics, and four prominent applications. We then explore prevalent methods for modeling diverse inputs, including structured NQG, unstructured NQG, and hybrid NQG. Despite the notable achievements of NQG models, several challenges remain, suggesting promising directions for future research.
@@
Multi-modal Question Generation. Current question generation tasks primarily concentrate on single-modal question generation, such as KBQG, TQG, and VQG. Nevertheless, in practical scenarios, the significance of multi-modal question generation is on the rise. This is particularly evident in the educational field, where numerous scientific questions require an understanding of both visual images and textual descriptions. To the best of our knowledge, multi-modal question generation remains in a very early stage, with few studies having been conducted in this field Wang and Baraniuk (2023), as detailed in Section 5. Given the advanced capabilities of vision-language pre-trained models like CLIP, a compelling research direction is to develop effective strategies for leveraging VL-PLMs in multi-modal question generation.
@@
Acknowledgments
This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and the National Natural Science Foundation of China (62322214, U23A20299, 62076245, 62072460, 62172424, 62276270). This work is supported by Public Computing Cloud, Renmin University of China. We gratefully acknowledge the support provided by the China Scholarship Council Scholarship Fund. We sincerely thank all reviewers for their valuable feedback.